solve differential equation

If - Dord, find y @ (5,12) dy + 2xy = 6xe*** (11 pts) dx 4. dx + x* ydy = 0
Answer #1

cus Integrating both lides, we get: - Lue 2x y = 3(2x-17* tc x = 5 and 4 =12 At 12 = 3 ( 2(5)-1) ² + c 12= 3 (9) * tc 12 = 3nlene, finally y = 3(2x-1) +9As per HomeworkLib policy, I have authority to do first question only. Thank you.